# Context-free graphs and their transition groups

### v2

@@ 5. Detecting torsion
Note that in the case of a finite union of context-free graphs $\{\Gamma^{(1)},\ldots,\Gamma^{(k)}\}$ the first fact of the proposition remains unchanged: whenever there are two self-similar end-cones in one of the graphs and a walk as in the above, the element has infinite order. While for the second fact, instead of considering $\delta$ and $N$, we need $\max_{i\in\{1,\ldots,k\}}\delta_{i}$ and $\max_{i\in\{1,\ldots,k\}}N_{i}$ where $\delta_{i}$ and $N_{i}$ are the corresponding parameters for $\Gamma^{(i)}$.
If the inverse graph $\Gamma$ is infinite, then, since there finitely many end-cone types and by the pigeonhole principle, we may deduce that there are at least two self-similar end-cones. This observation and Proposition 5.3 lead to the following fact.
@@
The transition group $\mathcal{G}(\Gamma^{(1)}\sqcup\ldots\sqcup\Gamma^{(k)})\in{\bf CF-TR}$ of a family of inverse graphs $\{\Gamma^{(1)},\ldots,\Gamma^{(k)}\}$ is infinite if and only if at least one graph among $\Gamma^{(1)},\ldots,\Gamma^{(k)}$ is infinite.
Note that given an inverse graph $\Gamma$ presented as an inverse PDA, it is undecidable to check whether $\mathcal{G}(\Gamma)$ is infinite or not. This because it is well-known that it is undecidable to tell whether a PDA is a FSA. Proposition 5.3 gives an alternative proof of the decidability of the torsion problem for the transition groups of context-free graphs. This result may also be derived from the same result for the class of co-context-free groups (see [17, Theorem 9]).
We have the following corollary.
@@
An infinite group $G=\mathcal{G}(\Gamma^{(1)},\ldots,\Gamma^{(k)})\in{\bf CF-TR}$ does not contain an infinite finitely generated torsion subgroup.
@@
This is a consequence of Proposition 4.14, Proposition 5.3 and Proposition 5.5, because any infinite finitely generated subgroup is still in ${\bf CF-TR}$ and so one of the connected components $\Gamma^{(i)}$ of the graph is infinite and it must have an element of infinite order. ∎
Remark 5.7.
By the previous corollary we have that no infinite torsion group is CF-TR, in particular, Grigorchuk group does not belong to such class. However, it is still open the case whether it is CoCF ([12]).
@@
With the above notation, for a given element $g\in\mathcal{G}(\Gamma^{(1)},\ldots,\Gamma^{(k)})$ there is an algorithm to check whether $o(g)$ is finite or not. In other words, the torsion problem is solvable for CF-TR groups.
Proof.
We deal with the case $\mathcal{G}(\Gamma)$, the general case is analog. By Proposition 5.3 if $o(g)$ is finite, then
(2) $\frac{o(g)}{\log_{2}(o(k))}<|g|\delta\gamma^{N|g|}$
Note that $\frac{o(g)}{\log_{2}(o(k))}$, as a function of $o(g)$, has strictly positive derivative for $o(g)>e$, thus for $o(g)\geq 3$ this function is monotonically increasing so there is a (computable) function $C(|g|):\mathbb{N}\to\mathbb{N}$ such that for any $3\leq o(g)<C(|g|)$, the bound (2) is satisfied. Therefore, a generic element $g$ has either finite order satisfying the bound $o(g)<C(|g|)$, or it has infinite order. Hence, to check whether $g$ has finite order it is enough to check if $g^{C(|g|)!}=1$ in $\mathcal{G}(\Gamma)$, or by simply checking whether $g^{k}=1$ for some $k\leq C(|g|)$. Since the word problem for ${\bf CoCF}$ is decidable (see [17, Proposition 7]), the former problem is also decidable. ∎
We know that for $F$ a free group of rank at least $2$, the group $F\times F$ is in CF-TR. Indeed, $F$ is in CF-TR simply by considering the Cayley graph on free generators and, by Proposition 4.15, the class is closed under direct products. This provides an example of a CF-TR group with unsolvable conjugacy problem (see e.g. [21, Theorem 9, Ch.3]) and unsolvable generalized word problem ([20, Theorem 4.6]).
Most of these groups have exponential growth. This happens if the tree which is quasi-isometric to $\Gamma$ has more than two ends. For context-free graphs, this is equivalent to having at least two self-similar end-cones $\Gamma(p,x_{0})\subseteq\Gamma(q,x_{0})$ such that $\Gamma(q,x_{0})$ has $k>1$ second level end-cones.
Proposition 5.9.
If $\Gamma$ is quasi-isometric to a tree with more than two ends, then $\mathcal{G}(\Gamma)$ has exponential growth.
Proof.
Let $V(D_{n}(x_{0}))=\{v\in V(\Gamma):\|v\|_{x_{0}}=n\}$ be the set of vertices $v$ which are at distance $n$ from the root $x_{0}$, and for each $v\in V(D_{n}(x_{0}))$ let us choose a word $w(v)\in A^{*}$ such that $x_{0}\smash{\stackrel{{\scriptstyle\text{\tiny{$w(v)$}}}}{{\relbar\joinrel\longrightarrow}}}v$ is a walk connecting $x_{0}$ with $v$. Let
$G_{n}=\{g_{v}\ \in\mathcal{G}(\Gamma):g_{v}\mbox{ is represented by }w(v)\mbox{ for each }v\in V(D_{n}(x_{0}))\}$
be the set of elements of $\mathcal{G}(\Gamma)$ corresponding to $w(v)$ for each $v\in V(D_{n}(x_{0}))$. Note that $|G_{n}|=|V(D_{n}(x_{0}))|$ since if $v_{1},v_{2}\in V(D_{n}(x_{0}))$ are two distinct vertices, then the corresponding elements $g_{v_{1}},g_{v_{2}}$ in $\mathcal{G}(\Gamma)$ are different since they acts differently on $x_{0}$. Moreover, each $g_{v}\in G_{n}$ is in the ball $B_{n}$ of radius $n$ in the Cayley graph of $\mathcal{G}(\Gamma)$. Thus, for each $n>0$ we have the exponential upper bound
$B_{n}\geq|G_{n}|=|V(D_{n}(x_{0}))|$
Remark 5.10.
If we are dealing with a finite union of graphs, we ask that at least one of them is quasi-isometric to a tree with more than two ends.
@@ 6. Poly-context-free groups and co-context-free transition groups
In this section, we try to shed some light on the first question of Problem 4.11. At the end of the section we also make some considerations on some conjectures regarding poly-context-free groups.
Throughout this section, we focus on groups whose word problem is the intersection of finitely many inverse-context-free languages, subsequently denoted as Poly-ICF. The following theorem shows that this class is contained in CF-TR.
@@
Let $G$ be a group whose word problem $WP(G;A)$ is the intersection of $k$ context-free languages accepted by rooted inverse graphs:
@@
The $A$-digraph
@@
with $\alpha=(x_{1},\ldots,x_{k})$ is clearly inverse since each $\Gamma_{i}$ is inverse, and since $WP(G;A)=L(\Gamma,\alpha)$, by Proposition 2.5 we deduce that $\Gamma$ is isomorphic to the Cayley graph $\operatorname{Cay}(G;A)$. We recall that for any $u\in A^{*}$ and $x\in V(\Gamma)$, $x\cdot u$ is the endpoint of the walk $x\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}(x\cdot u)$ in an inverse and complete $A$-digraph $\Gamma$. We clearly have $\alpha\cdot u=(x_{1}\cdot u,\ldots,x_{k}\cdot u)$, furthermore, it is straightforward to check that the following equality holds for all $u\in A^{*}$:
[…]
Now, since $\Gamma$ is a Cayley graph $(\Gamma,\alpha)\simeq(\Gamma,\alpha\cdot u)$, thus for all $u\in A^{*}$ we have $WP(G;A)=\bigcap_{i=1}^{k}L(\Gamma,x_{i}\cdot u)$. Hence, by recalling equation (1), we conclude that
(3) $WP(G;A)=\bigcap_{i=1}^{k}\left(\bigcap_{y\in V(\Gamma^{(i)})}L(\Gamma^{(i)},y)\right)=\bigcap_{i=1}^{k}\mathcal{L}(\Gamma^{(i)})=\mathcal{L}(\Gamma^{(1)},\ldots,\Gamma^{(k)})$
i.e., $G\simeq\mathcal{G}(\Gamma^{(1)}\sqcup\ldots\sqcup\Gamma^{(k)})$. ∎
An interesting question is whether there exists examples of CF-TR groups which are not poly-context-free. The next result tells us that such examples needs to be non-quasi-transitive (for a specific example, refer to Section 10). More precisely, it gives a characterization that offers a kind of converse to the previous theorem.
@@
Let $\{\Gamma^{(1)},\ldots,\Gamma^{(k)}\}$ be a collection of quasi-transitive context-free inverse $A$-digraphs. Then, $G=\mathcal{G}(\Gamma^{(1)}\sqcup\ldots\sqcup\Gamma^{(k)})$ is a group in CF-TR if and only if $G$ is virtually a finitely generated subgroup of the direct product of free groups. In particular, it belongs to the class of poly-context-free groups whose word problem is the intersection of finitely many languages accepted by rooted quasi-transitive context-free inverse graphs.
@@
Fix a generic inverse $A$-digraph $\Gamma^{(i)}$ of the family $\{\Gamma^{(1)},\ldots,\Gamma^{(k)}\}$. Since it is quasi-transitive, there is a finite complete set $F_{i}\subseteq V(\Gamma^{(i)})$ of representatives of the equivalence relation given by the orbits of $\operatorname{Aut}(\Gamma_{i})$. Since by Proposition 2.5 we have that for any $\varphi\in\operatorname{Aut}(\Gamma^{(i)})$, $L(\Gamma^{(i)},v)=L(\Gamma^{(i)},\varphi(v))$ we deduce:
@@
is the intersection of finitely many context-free languages accepted by quasi-transitive inverse $A$-digraphs, hence by [24, Theorem 3] this is equivalent for the group $G$ to be virtually a finitely generated subgroup of the direct product of free groups. Conversely, by [24, Proposition 12] if $G$ is virtually a finitely generated subgroup of the direct product of free groups, then the word problem $WP(G;A)=\bigcap_{i=1}^{k}L(\Gamma^{(i)},x_{i})$ is the intersection of finitely many inverse-context-free languages accepted by inverse graphs that are quasi-transitive. Hence, by Theorem 6.1 we have that the group $G\simeq\mathcal{G}(\Gamma^{(1)}\sqcup\ldots\sqcup\Gamma^{(k)})$ is the transition group of the disjoint union of finitely many inverse context-free $A$-digraphs that are quasi-transitive. ∎
The previous theorem offers another characterization of groups that are virtually finitely generated subgroups of the direct product of free groups. Indeed, in [24, Theorem 3.1] it is shown that such class is also characterized by the groups whose word problem is in Poly-QT-ICF, i.e., languages that are the intersection of finitely many languages accepted by inverse graphs that are quasi-transitive and quasi-isometric to a tree. Note that the example presented in subsection 4.1 is in Poly-QT-ICF since by a quick inspection of Fig. 1 shows that $\Omega$ is quasi-transitive.
Following [8] we give the following construction. For $c=\langle c_{0},\ldots,c_{s}\rangle\in\mathbb{Z}^{s+1}$ with $s\geq 1,c_{0},c_{s}\neq 0$ and $\mathrm{GCD}(c_{0},\ldots,c_{s})=1$, the group $G(c)$ is defined by the presentation $\langle a,b\mid\mathfrak{R}_{c}\rangle$, where $\mathfrak{R}_{c}=\{[b,b^{a^{i}}](i\in\mathbb{Z}),b^{c_{0}}(b^{a})^{c_{1}}\ldots(b^{a^{s}})^{c_{s}}\}$ via the notation $x^{y}:=y^{-1}xy$ and $[x,y]:=x^{-1}y^{-1}xy$. These groups are called $Gc$-groups and we say that a $Gc$-group is proper if it is not virtually abelian.
We end this section with some consideration on Brough Conjecture 2 where it is conjectured that a finitely generated poly-context-free group is solvable if and only if it is virtually abelian. In this direction we have the following proposition.
@@
By assumption the group $G$ is CF-TR then it is CoCF and so, since by [9, Proposition 4.13] a $Gc$-group is poly-context-free or co-context-free if and only if it is virtually abelian, we deduce that case (2) is not possible. Let us now analyze case (3) and assume that $G$ contains a subgroup $H$ with an infinite normal torsion subgroup $U$. Suppose that $U=\langle g_{i}:i\in\mathbb{N}\rangle$ and consider the subgroups $U_{n}=\langle g_{i}:i\leq n\rangle\leq U$. The subgroups $U_{n}$ are clearly finitely generated torsion subgroups of $G$ and thus, by Corollary 5.6, they are all finite. Clearly, the order of the subgroups $U_{n}$ becomes arbitrarily large since $U$ is infinite. ∎
@@
(1) $\implies$ (2) Since $\mathbb{Z}/n\mathbb{Z}$ and $\mathbb{Z}$ can be realized as transition groups of context-free graphs by Example 4.2, then Proposition 4.15 and Proposition 4.18 imply that finitely generated virtually abelian groups are also in the class of transition groups as well as poly-context-free by [9]. (2) $\implies$ (1) The converse implication follows by Proposition 6.3 and the fact that $\mathbb{Z}^{\infty}$ is not poly-context-free, as it contains free abelian subgroups of rank $k$ for all $k\in\mathbb{N}$, see [9, Lemma 4.6]. ∎
From the previous discussion we have the following natural questions.
@@
  • Is it possible to realize a solvable transition group associated to a context-free graph containing $\mathbb{Z}^{\infty}$ as a subgroup? In particular, is the group described in Definition 10.4 solvable?
@@
  • Is it true that a transition group associated to a context-free graph does not have arbitrarily large finite subgroups? By the argument at the end of the proof of Proposition 6.3 this is equivalent to not having an infinite torsion subgroup.
7. Perturbing a graph: a structural result
A natural question about these groups is how much small local changes in the graph affect the associated transition groups. In this section we investigate what happens to the transition group when we locally perturb a context-free graph $(\Gamma,x_{0})$. The idea is to consider a ball $D_{n}(x_{0})$ and to associate with the inverse graph $\Lambda_{n}=\Gamma\setminus D_{n}(x_{0})$ a group called the boundary group $\mathrsfs{B}_{n}$ at level $n$, and then to give a structural result connecting this group to the transition group $G=\mathcal{G}(\Gamma)$.
Definition 7.1 (Boundary group).
Consider the following subset of $A^{*}$
$\mathcal{O}_{n}=\{u\in A^{*}:\mbox{ if }x\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}x^{\prime}\mbox{ is a walk in }\Lambda_{n}\mbox{ then }x=x^{\prime}\}$
We have the following key lemma.
Lemma 7.2.
Proof.
Consider the infinite walk $x\smash{\stackrel{{\scriptstyle\text{\tiny{$u^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots$, and let us denote by $x_{i}^{(j)}$ the vertices of such walk with $x=x_{0}^{(0)}$ and $x_{i}^{(j)}$ for $0\leq i\leq|u|$ are all the vertices between $x_{0}^{(j)}$ and $x_{0}^{(j+1)}$ so that $x_{0}^{(j)}\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}x_{0}^{(j+1)}$ are the corresponding subwalks of $x\smash{\stackrel{{\scriptstyle\text{\tiny{$u^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots$. If $x\cdot u=x$ we are done, thus we may suppose $x\cdot u\neq x$. Note that every subwalk $x_{0}^{(j)}\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}x_{0}^{(j+1)}$ intersects $S_{n}(x_{0})$. Indeed, if $x_{0}^{(j)}\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}x_{0}^{(j+1)}$ were totally contained in $\Lambda_{n}$, then since $u\in\mathcal{O}_{n}$, we would have $x_{0}^{(j)}=x_{0}^{(j+1)}$, hence $x\cdot u=x$ when $j=0$, a contradiction. Now if we have that $x_{i}^{{(j)}}=x_{i}^{{(s)}}=v\in S_{n}(x_{0})$, for some $s\neq j$, then by the determinism of $\Gamma$ we have $x\cdot u^{k}=x$ for some $k$. Since the walk $x\smash{\stackrel{{\scriptstyle\text{\tiny{$u^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots$ is infinite and $S_{n}(x_{0})$ is a finite set. Either one of the two cases happens. To give the upper bound on $k$, we consider worst case scenario: each subwalk $x_{0}^{(j)}\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}x_{0}^{(j+1)}$ contains at most a vertex of $S_{n}(x_{0})$, therefore by the pigeonhole principle it is easy to see that condition $x_{i}^{{(j)}}=x_{i}^{{(s)}}$ occurs for some $j,s\leq\left|S_{n}(x_{0})\right|\,\left|u\right|$. ∎
Corollary 7.3.
For all $n\geq 1$, the boundary group $\mathrsfs{B}_{n}$ at level $n$ is a torsion normal subgroup of $G$ where each element $g$ has order $o(g)\leq\left|S_{n}(x_{0})\right|\,\left|g\right|$ (where ||$\cdot$|| ⋅ | is the group metric with respect to $A$).
In view of the structural result contained in Proposition 6.2 an interesting case is when the graph $\Lambda_{n}$ can be obtained by a quasi-transitive inverse graph by cutting out the disk $D_{n}(x_{0})$, more precisely we define the following class of inverse graphs.
Definition 7.4.
An inverse graph $(\Gamma,x_{0})$ is called locally-quasi-transitive if there is a quasi-transitive inverse graph $(\Theta,y_{0})$ and an integer $n\geq 1$ such that $\Gamma\setminus D_{n}(x_{0})=\Lambda_{n}=\Theta\setminus D_{n}(y_{0})$.
Lemma 7.5.
$\mathcal{O}_{n}=\bigcap_{p\in R}L(\Theta,p)=\mathcal{L}(\Theta)$
Proof.
The equality $\bigcap_{p\in R}L(\Theta,p)=\mathcal{L}(\Theta)$ follows by Proposition 2.5 as two vertices in the same orbits have the same associate language. Since $\bigcap_{p\in R}L(\Theta,p)=\mathcal{L}(\Theta)$ are words that label a circuit at every point, we clearly have the inclusion $\bigcap_{p\in R}L(\Theta,p)\subseteq\mathcal{O}_{n}$. Conversely, take a word $u\in\mathcal{O}_{n}$, for each $p_{i}\in R$ we claim that $u$ labels a circuit at $p_{i}$. Indeed, since the orbit $O_{i}$ is infinite, we may find a vertex $p_{i}^{\prime}\in O_{i}$ that is at sufficiently large distance from the disk $D_{n}(x_{0})$ such that $p_{i}^{\prime}\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}z$ is totally contained in $\Lambda_{n}$, hence by definition $z=p_{i}^{\prime}$ and therefore $u$ labels also a circuit at $p_{i}$. Therefore, $u\in\bigcap_{p\in R}L(\Theta,p)$ showing the other inclusion. ∎
In case the inverse graph $(\Gamma,x_{0})$ is locally-quasi-transitive where the associated quasi-transitive graph $\Theta$ has all the orbits that are infinite, we say that $(\Gamma,x_{0})$ is locally-quasi-transitive with infinite orbits. We are now in position to prove the following structural theorem.
Theorem 7.6.
Let $(\Gamma,x_{0})$ be a context-free inverse graph that is also locally-quasi-transitive with infinite orbits and let $G=\mathcal{G}(\Gamma)$ its transition group. Then, there is a short exact sequence
Proof.
Let $\Theta$ be the inverse graph which is quasi-transitive with infinite orbits, such that there exists an integer $n\geq 1$ with $\Gamma\setminus D_{n}(x_{0})=\Lambda_{n}=\Theta\setminus D_{n}(y_{0})$. Since there are finitely many end-cone types with frontier vertices contained inside the ball of radius $n$, we have that $\Gamma$ and $\Theta$ differ from finitely many end-cones up to end-isomorphism. Hence, $\Theta$ is also context-free. Therefore, by Proposition 6.2 we have that $H=\mathcal{G}(\Theta)$ is virtually a finitely generated subgroup of the direct product of free groups. Now, by Lemma 7.5 we have $\mathcal{O}_{n}=\mathcal{L}(\Theta)$, hence since $\mathcal{L}(\Gamma)\subseteq\mathcal{O}_{n}=\mathcal{L}(\Theta)$, there is a well-defined epimorphism $\psi:\mathcal{G}(\Gamma)\rightarrow\mathcal{G}(\Theta)$, with $\ker(\psi)=\pi(\mathcal{O}_{n})=\mathrsfs{B}_{n}$ ($\pi$ is the natural map from $A^{*}$ to the group $\mathcal{G}(\Gamma)$). From which we have the short exact sequence in the statement and the fact that $H\simeq G/_{\mathrsfs{B}_{n}}$. Moreover, by Corollary 7.3, we have that the boundary group $\mathrsfs{B}_{n}$ is a torsion subgroup of $\mathcal{G}(\Gamma)$ where the order of each element $g$ is bounded above by $o(g)\leq\left|S_{n}(x_{0})\right|\,\left|g\right|$. ∎
It is worth mentioning that in the settings of the previous proposition, the boundary group $\mathrsfs{B}_{n}$ is generated by a poly-context-free language. In Section 10 we will show a concrete example of a transition group associated to a context-free locally-quasi-transitive graph having boundary group which is infinite and whose elements have order two. This gives a negative answer to the second point of Question 6.5.

### v3

@@ 5. Detecting torsion
Note that in the case of a finite union of context-free graphs $\{\Gamma^{(1)},\ldots,\Gamma^{(k)}\}$, the first part of the proposition clearly remains valid. As for the second part, instead of considering the parameters $\delta$ and $N$, one must take $\max_{i\in\{1,\ldots,k\}}\delta_{i}$ and $\max_{i\in\{1,\ldots,k\}}N_{i},$ where $\delta_{i}$ and $N_{i}$ denote the respective parameters of $\Gamma^{(i)}$.
Question 5.6.
Is there a CF-TR group realizing the upper bound provided in the second point of Proposition 5.4?
As a consequence of the previous proposition we have:
@@
Let $\{\Gamma^{(1)},\ldots,\Gamma^{(k)}\}$ be a collection of context-free graphs, and let $G=\mathcal{G}\bigl(\Gamma^{(1)}\sqcup\cdots\sqcup\Gamma^{(k)}\bigr)\in\mathbf{CF\text{-}TR}$ be the associated transition group. Then, the following are equivalent:
  1. (1)
    $G$ contains a copy of $\mathbb{Z}$;
  2. (2)
    $G$ is infinite;
  3. (3)
    at least one of the graphs $\Gamma^{(1)},\ldots,\Gamma^{(k)}$ is infinite.
Proof.
$(1)\Rightarrow(2)$ is trivial. For $(2)\Rightarrow(3)$, note that if all $\Gamma^{(i)}$ are finite, then $G$ is finite. For $(3)\Rightarrow(1)$, assume without loss of generality that $\Gamma^{(1)}$ is infinite. Since there are finitely many end-cone types but infinitely many end-cones, the pigeonhole principle guarantees the existence of at least two nested end-cones. By Proposition 5.4, this produces an element of infinite order. ∎
Another consequence of the previous facts is the following.
@@
An infinite group $\mathcal{G}\bigl(\Gamma^{(1)}\sqcup\cdots\sqcup\Gamma^{(k)}\bigr)\in{\bf CF-TR}$ does not contain an infinite finitely generated torsion subgroup.
@@
This follows from the closure property under taking finitely generated subgroups, as stated in Proposition 4.17 and Proposition 5.7. ∎
It remains an open problem whether the Grigorchuk group belongs to the class CoCF (see [13]), and more generally, whether there are infinite torsion groups in the class CoCF. However, using the previous corollary, we can show the following.
Corollary 5.9.
No infinite torsion group can be CF-TR. In particular, the Grigorchuk group is not CF-TR.
Proposition 5.4 provides an alternative proof of the decidability of the torsion problem for the transition groups of context-free graphs by checking whether $g^{k}=\mathds{1}_{G}$ for some $k\leq e^{-N\delta}\,\delta^{N\delta|g|}\,\gamma^{\delta^{3}N^{2}|g|^{2}}.$ This decidability result can also be deduced from the corresponding result for the class of co-context-free groups (see [18, Theorem 9]).
@@
With the above notation, for a given element $g\in\mathcal{G}\bigl(\Gamma^{(1)}\sqcup\cdots\sqcup\Gamma^{(k)}\bigr)$ there is an algorithm to check whether $o(g)$ is finite or not. In other words, the torsion problem is solvable for CF-TR groups.
We know that for $F$ a free group of rank at least $2$, the group $F\times F$ is in CF-TR. Indeed, $F$ is in CF-TR simply by considering the Cayley graph on free generators and, by Proposition 4.18, the class is closed under direct products. This provides an example of a CF-TR group with unsolvable conjugacy problem (see e.g. [24, Theorem 9, Ch.3]) and unsolvable generalized word problem ([23, Theorem 4.6]).
@@ 6. Poly-context-free groups and co-context-free transition groups
In this section, we address the first question of Problem 4.13 and, at the end, discuss some conjectures on poly-context-free groups. We first focus on groups whose word problem is the intersection of finitely many inverse-context-free languages (denoted Poly-ICF), i.e., intersections of context-free languages accepted by graphs. The following theorem shows that this class is contained in CF-TR; it corresponds to the special case of intersections of deterministic context-free languages, which are contained in Co-CF.
@@
Let $G$ be a group whose word problem $WP(G;A)$ is the intersection of $k$ context-free languages accepted by rooted graphs:
@@
Let $(\Gamma,\alpha)=\bigtimes_{i=1}^{k}(\Gamma^{(i)},x_{i})$ be the $A$-graph defined as the product of the rooted graphs $(\Gamma^{(i)},x_{i})$. The vertex set of $\Gamma$ is the Cartesian product $V(\Gamma)=V(\Gamma^{(1)})\times\cdots\times V(\Gamma^{(k)})$, so each vertex is a $k$-tuple $(v_{1},\ldots,v_{k})$. A transition labeled by $a\in A$ acts componentwise:
@@
The root of $\Gamma$ is $\alpha=(x_{1},\ldots,x_{k})$. Since each $\Gamma^{(i)}$ is inverse, it is immediate that $\Gamma$ is also inverse. Moreover, since $WP(G;A)=L(\Gamma,\alpha)$, Proposition 2.5 implies that $\Gamma\simeq\operatorname{Cay}(G;A)$. Recall that for any $u\in A^{*}$ and $x\in V(\Gamma)$, $x\cdot u$ is the endpoint of the walk $x\xrightarrow{u}(x\cdot u)$ in a complete $A$-graph. Then $\alpha\cdot u=(x_{1}\cdot u,\ldots,x_{k}\cdot u)$ and $(\Gamma,\alpha\cdot u)=\bigtimes_{i=1}^{k}(\Gamma^{(i)},x_{i}\cdot u)$. Since $\Gamma$ is a Cayley graph, $(\Gamma,\alpha)\simeq(\Gamma,\alpha\cdot u)$ for all $u\in A^{*}$, and so $WP(G;A)=\bigcap_{i=1}^{k}L(\Gamma,x_{i}\cdot u)$ for all $u\in A^{*}$. Using (1), we obtain
$WP(G;A)=\bigcap_{i=1}^{k}\left(\bigcap_{y\in V(\Gamma^{(i)})}L(\Gamma^{(i)},y)\right)=\bigcap_{i=1}^{k}\mathcal{L}(\Gamma^{(i)})=\mathcal{L}(\Gamma^{(1)},\ldots,\Gamma^{(k)}),$
i.e., $G\simeq\mathcal{G}(\Gamma^{(1)}\sqcup\cdots\sqcup\Gamma^{(k)})$. ∎
An interesting question is whether there exist examples of CF-TR groups that are not poly-context-free. The next result shows that any such example must be non-quasi-transitive (for a specific example, see Section 10). More precisely, it provides a characterization that can be seen as a kind of converse to the previous result.
@@
Let $\{\Gamma^{(1)},\ldots,\Gamma^{(k)}\}$ be a collection of quasi-transitive context-free $A$-graphs. Then, $G=\mathcal{G}(\Gamma^{(1)}\sqcup\ldots\sqcup\Gamma^{(k)})$ is a group in CF-TR if and only if $G$ is virtually a finitely generated subgroup of the direct product of free groups. In particular, it belongs to the class of poly-context-free groups whose word problem is the intersection of finitely many languages accepted by rooted quasi-transitive context-free graphs.
@@
Fix a generic $A$-graph $\Gamma^{(i)}$ of the family $\{\Gamma^{(1)},\ldots,\Gamma^{(k)}\}$. Since it is quasi-transitive, there is a finite complete set $F_{i}\subseteq V(\Gamma^{(i)})$ of representatives of the equivalence relation given by the orbits of $\operatorname{Aut}(\Gamma_{i})$. Since by Proposition 2.5 we have that for any $\varphi\in\operatorname{Aut}(\Gamma^{(i)})$, $L(\Gamma^{(i)},v)=L(\Gamma^{(i)},\varphi(v))$ we deduce:
@@
is the intersection of finitely many context-free languages accepted by quasi-transitive $A$-graphs, hence by [27, Theorem 3] this is equivalent for the group $G$ to be virtually a finitely generated subgroup of the direct product of free groups. Conversely, by [27, Proposition 12] if $G$ is virtually a finitely generated subgroup of the direct product of free groups, then the word problem $WP(G;A)=\bigcap_{i=1}^{k}L(\Gamma^{(i)},x_{i})$ is the intersection of finitely many inverse-context-free languages accepted by graphs that are quasi-transitive. Hence, by Theorem 6.1 we have that the group $G\simeq\mathcal{G}(\Gamma^{(1)}\sqcup\ldots\sqcup\Gamma^{(k)})$ is the transition group of the disjoint union of finitely many context-free $A$-graphs that are quasi-transitive. ∎
The previous theorem provides another characterization of groups that are virtually finitely generated subgroups of direct products of free groups. In particular, [27, Theorem 3.1] shows that this class coincides with groups whose word problem lies in Poly-QT-ICF, i.e., intersections of finitely many languages accepted by graphs that are quasi-transitive and quasi-isometric to a tree. Note that the example in subsection 4.2 belongs to Poly-QT-ICF, since a quick inspection of Fig. 1 shows that $\Omega$ is quasi-transitive.
Following [9], for $c=\langle c_{0},\ldots,c_{s}\rangle\in\mathbb{Z}^{s+1}$ with $s\geq 1$, $c_{0},c_{s}\neq 0$, and $\mathrm{GCD}(c_{0},\ldots,c_{s})=1$, the group $G(c)$ is defined by
$G(c)=\langle a,b\mid\mathfrak{R}_{c}\rangle,\quad\mathfrak{R}_{c}=\{[b,b^{a^{i}}]\ (i\in\mathbb{Z}),\ b^{c_{0}}(b^{a})^{c_{1}}\cdots(b^{a^{s}})^{c_{s}}\},$
using the usual notation $x^{y}:=y^{-1}xy$ and $x,y]:=x^{-1}y^{-1}xy$. These are called $Gc$-groups, and a $Gc$-group is proper if it is not virtually abelian. We conclude this section with some remarks on Brough Conjecture 2, which asserts that a finitely generated poly-context-free group is solvable if and only if it is virtually abelian. In this direction, we have the following proposition.
@@
By assumption the group $G$ is CF-TR then it is CoCF and so, since by [10, Proposition 4.13] a $Gc$-group is poly-context-free or co-context-free if and only if it is virtually abelian, we deduce that case (2) is not possible. Let us now analyze case (3) and assume that $G$ contains a subgroup $H$ with an infinite normal torsion subgroup $U$. Suppose that $U=\langle g_{i}:i\in\mathbb{N}\rangle$ and consider the subgroups $U_{n}=\langle g_{i}:i\leq n\rangle\leq U$. The subgroups $U_{n}$ are clearly finitely generated torsion subgroups of $G$ and thus, by Corollary 5.8, they are all finite. Clearly, the order of the subgroups $U_{n}$ becomes arbitrarily large since $U$ is infinite. ∎
@@
(1) $\implies$ (2) Since $\mathbb{Z}/n\mathbb{Z}$ and $\mathbb{Z}$ can be realized as transition groups of context-free graphs by Example 4.3, then Proposition 4.18 and Proposition 4.21 imply that finitely generated virtually abelian groups are also in the class of transition groups as well as poly-context-free by [10]. (2) $\implies$ (1) The converse implication follows by Proposition 6.3 and the fact that $\mathbb{Z}^{\infty}$ is not poly-context-free, as it contains free abelian subgroups of rank $k$ for all $k\in\mathbb{N}$, see [10, Lemma 4.6]. ∎
The previous discussion raises the following natural questions, which will be addressed in the following sections.
@@
  • Is it possible to realize a solvable transition group associated to a context-free graph containing $\mathbb{Z}^{\infty}$ as a subgroup? See Propositon 10.5.
@@
  • Is it true that a transition group associated to a context-free graph does not have arbitrarily large finite subgroups? By the argument at the end of the proof of Proposition 6.3 this is equivalent to not having an infinite torsion subgroup. See the Section 8 and Proposition 10.6.